Grade 11Mechanics

A wooden block of mass 10 kg is suspended with a long string. A bullet of mass 50 g strikes the block and gets embedded in it. As a result, block swings to a maximum height of 10 cm. Find the initial velocity of the bullet.

archi
10 Years agoGrade 11

1 Answer

Amogh Gajare
ApprovedApproved Tutor Answer10 Years ago
Hello!
Principles used in the problem
Principle of conservation of momentum & Work Energy theorem.
 
Given information -
Mass of bullet = 50 g = 0.05 kg
Mass of block = 10 kg
 
Solution -
To find Velocity ‘v’ of bullet,
By principle of conservation of linear momentum,
m_1*v_1 = m_2*v_2
Initially, mass in block was just 10 kg.But after bullet got embeeded into it, bullet’s weight got added.
So,
0.05*v_1 = (10 + 0.05)*v_2…..................Equation 1
By work energy theorem,
Work done = change in kinetic energy
- mgh = K_f - K_i.............as reference level is taken above, it gives a negative sign to work)
(K\rightarrowKinetic energy)
- mgh = \frac{1}{2}mv_1^2 - \frac{1}{2}mv_2^2
As block was initially at rest, V1 = 0
So
- mgh = - \frac{1}{2}mv_2^2             
Substituting values and calculating we get,
v_2 = \sqrt2…............................................................Equation 2
From Equation 1 & 2 we get,
0.05*v_1 = 10.05*\sqrt2
On calculation,
v_1 = 201\sqrt2 m/s
If any further doubts, feel free to email me at -amoghsgajare @gmail.com